44. One type of jet engine has a 0.0001 probability of failure while in flight. For a jet that has four of these engines, what is the probability of at least two of them failing? Geometric distribution, show work
Answer #1

This is binomial distribution and not geometric

X follow binomial with n = 4 , p = 0.0001

P(X >= 2) = 1 - (P(X = 0) +P(X = 1))

= 1 - (q^n + np q^(n-1))

= 1 - ( (1-0.0001)^4 + 4 * 0.0001 * (1 - 0.0001)^3)

= 5.99920001e-8